Put the equation y=4x^2+7x+6 in vertex form and state the values of a, h, and k. Enter the exact answers.
Vertex form= a(x-h)^2+k

Answer by stanbon(75887)

Put the equation y=4x^2+7x+6 in vertex form and state the values of a, h, and k. Enter the exact answers.
Vertex form= a(x-h)^2+k
------
y-6 = 4(x^2+(7/4)x)
-----
4(x^2+(7/4)x+(7/8)^2) = y - 6 +4(7/8)^2
-----
4(x+(7/8))^2 = y - 6 + 49/16
------
4(x+(7/8))^2 = y-(47/16)
-----
y = 4(x+(7/8))^2 + (47/16)
-------
a = 4
h = -7/8
k = 47/16
------------
Cheers,
Stan H.
-------------